OpenStudy (anonymous):

The half-life of cobalt 60 is 5.3 years. What fraction of a sample remains after 21.2 years? A. one half B. one quarter C. one eighth D. one sixteenth

OpenStudy (anonymous):

\[\huge(\frac{1}{2})^{\frac{21.2}{5.3}}\]

OpenStudy (anonymous):

21.2 / 5.3 = 4 times. thus the answer is (1/2)^ 4 1/16